Find the magnitude and direction of v = 5i + 4j
Answer by stanbon(75887)

Find the magnitude and direction of v = 5i + 4j
----
magnitude = sqrt(5^2 + 4^2] = sqrt(41)
---
direction = ?
tan(x) = 4/5
x = tan^-1(4/5) = 38.66 degrees
------------------------------------
Cheers,
Stan H.